# LAMMPS-KOKKOS: Performance Portable Molecular Dynamics Across Exascale Architectures

### v1

@@ 4.1. Simple Pairwise Interatomic Potentials (Lennard-Jones)
There are two neighbor list styles: “half” and “full”. Using “half” neighbor lists exploits Newton’s third law, avoiding the duplicate computation of symmetric forces between two atoms. With “full” neighbor lists the force of an atom i onto an atom k is computed separately from the force of the atom k onto the atom i. For simple pairwise potentials, using a “half” neighbor list can cause a data write conflict between two threads when writing to the atom force array, which needs to be handled via a deconflicting approach, namely thread atomic operations or data duplication. Data deconflicting is handled transparently by Kokkos ScatterView objects. Alternatively, one can use a “full” neighbor list, which does not require atomics but duplicates work (every atom’s force is computed twice). Separately, LAMMPS has the option to use Newton’s third law for ghost atoms (copies of atoms owned by other MPI ranks or on the other side of a periodic boundary), which reduces computation but increases the amount of communication required. For simple pair-wise potentials, typically a full neighbor list and newton off is better for GPUs, while a half list and newton on is better for CPUs (in particular in the common case when using one MPI rank per core, and not multi-threading).
Which neighbor list style to use does not have a one-size-fits-all answer. It highly depends on the hardware architecture, the specific pair style, and the cutoff distance for the force computation. 2 demonstrates the impact of differing combinations of neighbor list setting combinations for the LJ potential for two different GPU architectures, NVIDIA H100 (red curves/squares) and AMD MI250X (blue curves/squares). Generally speaking, the more compute intensive a pair style is the more likely it is that half neighbor lists are the right choice. Furthermore, on some architectures, such as NVIDIA GPUs, the atomic throughput is high enough that the overhead of atomics can be lower than the cost of the redundant computation performed in full neighbor lists.
@@ 4.2. ReaxFF
In addition to the basic interatomic interactions described in 4.1 above, molecular dynamics codes also support potentials that explicitly model bonds between atoms. The addition of bonded states allows the simulation of entirely new categories of materials, such as biomolecules, catalysts, and polymers. LAMMPS contains a number of ways to model bonds in materials, and especially important amongst them is ReaxFF, or the Reactive Force Field (Van Duin et al., 2001). The term “reactive” in this context refers to the ability to not only model basic (static) bond parameters such as bond distances, angles, and energies, but also the dynamic formation and dissociation of bonds during a simulation. This additional capability allows for the simulation of complex chemical processes and the study of reaction pathways and kinetics (Senftle et al., 2016).
@@
The last step was unrelated to the programming model, but was important for exascale preparedness. This was proactively considering sources of 32-bit integer overflow seen in large simulations and replacing variables with 64 bit integers as appropriate. We discuss this in B [^1] See the following GitHub pull requests: https://github.com/lammps/lammps/pull/4207 and https://github.com/lammps/lammps/pull/4318..
The following section will focus on four kernels to flesh out the above discussion. First, to motivate work pre-processing to reduce divergence, we separately discuss the bond order neighbor list kernel and the four-body force kernel. Second, to motivate increasing parallelism to improve performance, we discuss the kernel that builds the sparse matrix in the charge equilibration step. Third, we briefly discuss fusing iterative sparse matrix solves to promote matrix reuse. Last, to motivate exascale preparedness, we are going to briefly discuss targeted usage of 64-bit integers in the context of the sparse matrix implementation. These kernels will be explored using a key LAMMPS ReaxFF benchmark, a short simulation of the molecular crystal Hexanitrostilbene (HNS).
@@ 4.2.1. ReaxFF Optimization: Reducing Divergence via Pre-Processing Kernels
The pre-processing phases determine the quads of indices $(i,j,k,l)$ that obey the aforementioned constraints and saves them to a Kokkos View of int4 data types. The first pre-processing kernel counts the total number of quads and stores the per-atom-$i$ count, the Kokkos View is resized if necessary, and the second pre-processing kernel stores the quads. The View of quads is populated using a thread-safe global queue. With this process all quads for an atom $i$ are guaranteed to be contiguous.
The updated computation kernel is now fully convergent because it only acts on quads obeying the requisite constraints. In addition, we can further improve performance by parallelizing over quads instead of atoms. Consecutive work-items will in general reuse some subset $i$, $j$, $k$, and $l$ values because all quads for atom $i$ are contiguous. This promotes cache reuse, boosting performance.
@@ 4.2.2. Exploiting hierarchical parallelism
This exposure was accomplished with Kokkos team hierarchical parallelism. Each Kokkos “thread” corresponds to one row. Individual matrix elements in a row are assigned to “vector” lanes, which map to GPU threads. Because per-row work is distributed over contiguous threads, convergent memory patterns are restored. This improves the performance of counting the number of non-zero elements per row (hierarchical reduction) as well as computing and slotting the values of the non-zero elements (hierarchical scan) into the modified CSR matrix data structures.
@@ 4.2.3. Kernel Fusion
The charge equilibration process is implemented as solving two linear systems to a fixed relative residual tolerance. The core kernel here is the sparse matrix-vector operation, which is bandwidth bound. The sparse matrix structure is the largest data structure. It is also identical between the two solves. By fusing the solves, performance improves by reusing the matrix load. We note this optimization was implemented by AMD for KOKKOS and was originally in the OpenMP version of ReaxFF (Aktulga et al., 2019).
@@ 4.3. SNAP
Significant effort has gone into optimizing both the SNAP machine learning algorithms and Kokkos implementation (Trott et al., 2014; Gayatri et al., 2020; Nguyen-Cong et al., 2021; LAMMPS, 2024). We describe some key points below.
In contrast to Lennard-Jones and ReaxFF, SNAP encodes the atomic neighborhood of a given atom $i$ in a more general hyperspherical decomposition. The hypherspherical basis on its own encodes two-body interactions. The decomposition for each (atom, neighbor) pair is encoded in Wigner U-matrices and summed. However, it is not the U-matrices but appropriate triple products thereof that are basis independent, and can encode up to four-body interactions.
@@
The Wigner U-matrices are the $\bf u_{j}$, which are a set of matrices of rank $2j+1$. The $j$ are positive integers and half-integers. The $\bf u_{j}$ are defined recursively; the right-hand side of 3 denotes this as a recursion relation where $\mathcal{F}$ is a linear operator mapping adjacent elements of ${\bf u}_{j-1/2}$ to elements of ${\bf u}_{j}$. The sum of these matrices over all neighbors $k$ of an atom $i$ gives the full description of the atomic neighborhood, ${\bf U}_{j}$.
@@
where $\otimes^{j}_{j_{1},j_{2}}$ denotes an $\mathcal{O}(j^{4})$ operation and :: represents an element-wise scalar product of matrices ($\mathcal{O}(j^{2})$). Group-theoretical symmetries allow us to constrain the computation to $0\leq j_{2}\leq j_{1}\leq j\leq J$, significantly reducing the required work and storage.
@@
Computing the SNAP force reduces to four steps which has a one-to-one mapping with the software implementation of the SNAP force evaluation. The initial, non-Kokkos CPU implementation proceeds as follows. There is an outer loop over all atoms. Within this outer loop, there are four subroutines corresponding to four steps of evaluating the SNAP force.
@@
    ComputeUi. Compute per-(atom, neighbor) pair Wigner $\bf u$ matrices, which are a representation of hyperpherical harmonics efficiently computed via a linear recursion relation, and accumulate them into the per-atom $\bf U$ matrices. This itself has an “intermediate” loop over neighbors and an inner-most loop over quantum numbers with a serial dependency.
@@
We lost the ability to reuse intermediate data structures across individual atom calculations when we switched to a parallel implementation. Each data structure needed a new “atom index” degree of freedom. This required a careful consideration of data layouts because, in the most general case, data locality and cache reuse is not guaranteed (Gayatri et al., 2020).
@@ 4.3.2. SNAP ComputeYi Implementation
ComputeYi is memory bound; computing each element of ${\bf Z}^{j}_{j_{1},j_{2}}$ requires weighted complex dot products of subsets of pairs of $\bf U$ matrices. By keeping temporal locality in computing $(j_{1},j_{2})$ pairs of $\bf Z$} components, we get cache reuse of ${\bf U}_{j_{1}}$ and ${\bf U}_{j_{2}}$. These can reside well in the lowest-level caches. Similarly, by preserving temporal locality in computing all components of ${\bf Z}^{j}_{j_{1},j_{2}}$; alternatively, locality in computing all of ${\bf Y}_{j}$, we can achieve cache reuse for the full set of $\bf U$ for a given atom.
@@
While we can transpose the data layout on the GPU, we cannot naïvely transpose the loop order and achieve locality. If, at fixed flattened $\bf Z$} index, we loop over all atom indices, the relevant components of $\bf U$ for the first atom falls out of lower level caches after traversing all atoms. The solution to this is a 3-d tiled traversal to get good data access patterns and locality.
@@
The ideal value of $v$ is architecture-dependent, with a non-trivial sensitivity to the granularity of work and to cache sizes. An intuitive lower bound is the warp (or equivalent) granularity, $v=32$ on NVIDIA GPUs and $v=64$ on AMD GPUs. The reasonable upper bound can be inferred by experiment but connects to L1 cache size. It was found in practice that the ideal values for $v$ are 32 on NVIDIA GPUs and 16 on Intel GPUs. Perhaps nonintuitively, the ideal value on AMD GPUs was seen to be v=32v=32.222These values are encoded at https://github.com/lammps/lammps/blob/4246fab5/src/KOKKOS/kokkos_type.h#L1379-L1387 Kokkos enables this explicit experimentation and tuning which is important for maximizing performance. A deeper discussion of the specific role of cache size can be found in 4.4.
@@ 4.3.3. ComputeUi and ComputeDuidrj
Further optimizations are possible on the GPU. This is based on the critical observation that recursive polynomial evaluation in inherently compute bound. The only inputs are the base case and all components of $\bf u$ follow from that.
@@
This approach carries over directly to ComputeDuidrj except we now stage the matrices $\bf U$} and the derivative $\bf\vec{d}u$}. In place to the atomic accumulation into $\bf U$} we can inline the accumulation into the force $\bf\vec{F}$—which is a kernel fusion with ComputeDeidrj. We denote this fused kernel ComputeFusedDeidrj.
@@ 4.3.4. Work Batching and Instruction Level Parallelism
We have already seen work batching in ReaxFF when the the two sparse linear system solves were fused, described in 4.2.3. In the sparse matrix-vector operation, the units of independent work are loading components of the right-hand-side vector, multiplying by the same matrix element, and accumulating into independent output vectors. Our limiter is loading the matrix elements. We can simultaneously hide the latency of loading both vectors behind the latency of loading the matrix elements. This is the origin of the performance improvement.
@@
We followed this same roadmap for implementing batching in the SNAP kernel. Our initial optimization efforts focused on NVIDIA H100. Limiters were identified using NVIDIA Nsight Compute. However, the idea of work batching to shift limiters and improve ILP is not specific to any one architecture. By introducing the batch size as a compile-time tunable parameter, we could find performance improvements on AMD MI250X and MI300A with otherwise identical code.
@@
We list the performance improvements for all three kernels in 2. As noted before, these optimizations were driven by a quantitative analysis of limiters on NVIDIA H100. We could then empirically re-tune the batching factors on AMD MI300A and see if we achieved a speed-up. In the “worst case”, a batching factor of $1$ would restore the original performance.
@@ 4.4. Cache Size Performance Impact
202040406060808010010012012014014016016018018020020022022024024000.50.511Shared memory carveout [KiB]Perf. relative to “default” carveoutComputeUiComputeYiComputeFusedDeidrjAllPairComputeLJCut
@@
One of the major performance limiters for molecular dynamics applications on CPUs is cache availability. This is true for GPU architectures as well. Generally speaking, GPUs have two sets of L1-like cache: a hardware and a software managed one. The latter is referred to as shared memory in CUDA, local data share in HIP, and local memory in SYCL.
The characteristics of these caches are one of the major differences of the GPU architectures used on exascale class systems (see 1). The sizes of the L1 cache and shared memory are fixed on AMD MI250X and MI300A, as well as on Intel PVC, because they are discrete units. On the other hand, modern NVIDIA GPUs have a unified cache where the L1 and shared memory capacity can be dynamically shifted.
The ability to shift between L1 and shared memory on NVIDIA GPUs enables us to design an experiment to investigate the impact of L1 or shared memory size on kernel performance leveraging what CUDA calls the shared memory carveout. This is the percentage of the unified cache reserved for shared memory. With kernels which do not use shared memory but benefit from L1 cache, we expect performance to decrease with an increasing ratio of the combined cache being reserved for shared memory. For kernels which use shared memory and do not rely on automatic caching in L1 the opposite is expected, and for ones which leverage both L1 and shared memory the best performance should be achieved with some intermediate carveout value.
Indeed Kokkos has a built-in heuristic to optimize the carveout value depending on the characteristics of the kernel launched. For this study, we overwrote that heuristic and simply forced a specific carveout value to see its impact on four top kernels: the primary force computation kernel for the Lennard Jones potential PairComputeLJCut, and the three most time consuming kernels in SNAP, ComputeUi, ComputeYi, and ComputeFusedDeidrj. Kernel runtimes were measured using NVIDIA Nsight Systems. We also looked at the top four kernels for the ReaxFF but did not see a significant dependence on the cache carveout (under 10%).
The effect of manually setting the cache carveout on NVIDIA H100-HBM3-SXM is shown in 3. The x-axis is the shared memory carveout and the y-axis is the performance of each kernel normalized by the performance from running at the default carveout.
The force kernel of Lennard Jones does not use shared memory, but does see significant benefits from large L1 cache sizes. At the maximum carveout for shared memory, which leaves only 32kB for L1, performance drops by about 55%; alternatively, increasing the L1 cache size from 32kB to 224kB increases performance by 85%. The ComputeYi kernel shows a similar behavior; in this case the difference comes from reuse of $\mathbf{U}_{j}$ matrices instead of neighbor atom coordinates.

### v2

@@ 4.1. Simple Pairwise Interatomic Potentials (Lennard-Jones)
There are two neighbor list styles: “half” and “full”. Using “half” neighbor lists exploits Newton’s third law, avoiding the duplicate computation of symmetric forces between two atoms. With “full” neighbor lists the force of an atom i onto an atom k is computed separately from the force of the atom k onto the atom i. For simple pairwise potentials, using a “half” neighbor list can cause a data write conflict between two threads when writing to the atom force array, which needs to be handled via a deconflicting approach, namely thread atomic operations or data duplication. Data deconflicting is handled transparently by Kokkos ScatterView objects. Alternatively, one can use a “full” neighbor list, which does not require atomics but duplicates work (every atom’s force is computed twice). Separately, LAMMPS has the option to use Newton’s third law for ghost atoms (copies of atoms owned by other MPI ranks or on the other side of a periodic boundary), which reduces computation but increases the amount of communication required. For simple pairwise potentials, typically a full neighbor list and newton off is better for GPUs, while a half list and newton on is better for CPUs (in particular in the common case when using one MPI rank per core, and not multi-threading).
Which neighbor list style to use does not have a one-size-fits-all answer. It highly depends on the hardware architecture, the specific pair style, and the cutoff distance for the force computation. Figure 2 demonstrates the impact of differing combinations of neighbor list settings for the LJ potential for two different GPU architectures, NVIDIA H100 (red curves/squares) and AMD MI250X (blue curves/squares). Generally speaking, the more compute intensive a pair style is the more likely it is that half neighbor lists are the right choice. Furthermore, on some architectures, such as NVIDIA GPUs, the atomic throughput is high enough that the overhead of atomics can be lower than the cost of the redundant computation performed in full neighbor lists.
@@ 4.2. ReaxFF
In addition to the basic interatomic interactions described in section 4.1 above, molecular dynamics codes also support potentials that explicitly model bonds between atoms. The addition of bonded states allows the simulation of entirely new categories of materials, such as biomolecules, catalysts, and polymers. LAMMPS contains a number of ways to model bonds in materials, and especially important amongst them is ReaxFF, or the Reactive Force Field (Van Duin et al., 2001). The term “reactive” in this context refers to the ability to not only model basic (static) bond parameters such as bond distances, angles, and energies, but also the dynamic formation and dissociation of bonds during a simulation. This additional capability allows for the simulation of complex chemical processes and the study of reaction pathways and kinetics (Senftle et al., 2016).
@@
The last step was unrelated to the programming model, but was important for exascale preparedness. This was proactively considering sources of 32-bit integer overflow seen in large simulations and replacing variables with 64-bit integers as appropriate. We discuss this in section appendix B [^1] See the following GitHub pull requests: https://github.com/lammps/lammps/pull/4207 and https://github.com/lammps/lammps/pull/4318..
The following section will focus on four kernels to flesh out the above discussion. First, to motivate work pre-processing to reduce divergence, we separately discuss the bond order neighbor list kernel and the four-body force kernel. Second, to motivate increasing parallelism to improve performance, we discuss the kernel that builds the sparse matrix in the charge equilibration step. Third, we briefly discuss fusing iterative sparse matrix solves to promote matrix reuse. These kernels will be explored using a key LAMMPS ReaxFF benchmark, a short simulation of the molecular crystal Hexanitrostilbene (HNS).
@@ 4.2.1. ReaxFF Optimization: Reducing Divergence via Pre-Processing Kernels
The pre-processing phases determine the quads of indices $(i,j,k,l)$ that obey the aforementioned constraints and saves them to a Kokkos View of int4 data types. The first pre-processing kernel counts the total number of quads and stores the per-atom-$i$ count, the Kokkos View is resized if necessary, and the second pre-processing kernel stores the quads. The View of quads is populated using a thread-safe global queue. With this process, all quads for an atom $i$ are guaranteed to be contiguous.
The updated computation kernel is now fully convergent because it only acts on quads obeying the requisite constraints. In addition, we can further improve performance by parallelizing over quads instead of atoms. Consecutive work-items will in general reuse some subset of $i$, $j$, $k$, and $l$ values because all quads for atom $i$ are contiguous. This promotes cache reuse, boosting performance.
@@ 4.2.2. Exploiting hierarchical parallelism
This exposure was accomplished with Kokkos team hierarchical parallelism. Each Kokkos “thread” corresponds to one row. Individual matrix elements in a row are assigned to “vector” lanes, which map to GPU threads. Because per-row work is distributed over contiguous threads, convergent memory patterns are restored. This improves the performance of counting the number of non-zero elements per row (hierarchical reduction) as well as computing and slotting the values of the non-zero elements (hierarchical scan) into the modified CSR matrix data structures. This optimization was found to be unprofitable on the CPU, so the code was bifurcated to allow the original version to run on CPUs, as described in 3.3.
@@ 4.2.3. Kernel Fusion
The charge equilibration process is implemented as solving two linear systems to a fixed relative residual tolerance. The core kernel here is the sparse matrix-vector operation, which is bandwidth bound. The sparse matrix structure is the largest data structure. It is also identical between the two solves. By fusing the solves, performance improves by reusing the matrix load. We note this optimization was implemented by AMD for the Kokkos version [^2] See https://github.com/lammps/lammps/pull/3147. and was originally in the OpenMP version of ReaxFF (Aktulga et al., 2019).
@@ 4.3. SNAP
Significant effort has gone into optimizing both the SNAP machine learning algorithms and Kokkos implementation (Trott et al., 2014; Gayatri et al., 2020; Nguyen-Cong et al., 2021; LAMMPS, 2024; Lasa et al., 2024). In particular, we note that Figure 6 of Ref. (Lasa et al., 2024) provides a partial history of the performance of SNAP over time, and shows that our optimizations for NVIDIA GPUs also gave speedup on AMD MI250X hardware. We describe some key points below.
In contrast to Lennard-Jones and ReaxFF, SNAP encodes the atomic neighborhood of a given atom $i$ in a more general hyperspherical decomposition. The hyperspherical basis on its own encodes two-body interactions. The decomposition for each (atom, neighbor) pair is encoded in Wigner U-matrices and summed. However, it is not the U-matrices but appropriate triple products thereof that are basis-independent, and can encode up to four-body interactions.
@@
The Wigner U-matrices are $\bf u_{j}$, which are a set of matrices of rank $2j+1$. The $j$ are positive integers and half-integers. The $\bf u_{j}$ are defined recursively; the right-hand side of equation 2 denotes this as a recursion relation where $\mathcal{F}$ is a linear operator mapping adjacent elements of ${\bf u}_{j-1/2}$ to elements of ${\bf u}_{j}$. The sum of these matrices over all neighbors $k$ of an atom $i$ gives the full description of the atomic neighborhood, ${\bf U}_{j}$.
@@
where $\otimes^{j}_{j_{1},j_{2}}$ denotes an $\mathcal{O}(j^{4})$ operation and “::” represents an element-wise scalar product of matrices ($\mathcal{O}(j^{2})$). Group-theoretical symmetries allow us to constrain the computation to $0\leq j_{2}\leq j_{1}\leq j\leq J$, significantly reducing the required work and storage.
@@
Computing the SNAP force reduces to four steps that have a one-to-one mapping with the software implementation of the SNAP force evaluation. The initial, non-Kokkos CPU implementation proceeds as follows. There is an outer loop over all atoms. Within this outer loop, there are four subroutines corresponding to four steps of evaluating the SNAP force.
@@
    ComputeUi. Compute per-(atom, neighbor) pair Wigner $\bf u$ matrices, which are a representation of hyperspherical harmonics efficiently computed via a linear recursion relation, and accumulate them into the per-atom $\bf U$ matrices. This itself has an “intermediate” loop over neighbors and an innermost loop over quantum numbers with a serial dependency.
@@
We lost the ability to reuse intermediate data structures across individual atom calculations when we switched to a parallel implementation. Each data structure needed a new “atom index” degree of freedom. This required a careful consideration of data layouts because, in the most general case, data locality and cache reuse are not guaranteed (Gayatri et al., 2020).
@@ 4.3.2. SNAP ComputeYi Implementation
The ComputeYi kernel requires computing each element of ${\bf Z}^{j}_{j_{1},j_{2}}$, which is the computation of weighted complex dot products of subsets of pairs of $\bf U$ matrices. This has a low computational intensity and trends memory bound. By keeping temporal locality in computing $(j_{1},j_{2})$ pairs of $\bf Z$} components, we get cache reuse of ${\bf U}_{j_{1}}$ and ${\bf U}_{j_{2}}$. These can reside well in the lowest-level caches. Similarly, by preserving temporal locality in computing all components of ${\bf Z}^{j}_{j_{1},j_{2}}$; alternatively, locality in computing all of ${\bf Y}_{j}$, we can achieve cache reuse for the full set of $\bf U$ for a given atom.
@@
While we can transpose the data layout on the GPU, we cannot naïvely transpose the loop order and achieve locality. If, at fixed flattened $\bf Z$} index, we loop over all atom indices, the relevant components of $\bf U$ for the first atom falls out of lower-level caches after traversing all atoms. The solution to this is a 3-d tiled traversal to get good data access patterns and locality, as inspired by the Cabana framework (Slattery et al., 2022).
@@
The ideal value of $v$ is architecture-dependent, with a non-trivial sensitivity to the granularity of work and to cache sizes. An intuitive lower bound is the warp (or equivalent) granularity, $v=32$ on NVIDIA GPUs and $v=64$ on AMD GPUs. The reasonable upper bound can be inferred by experiment but connects to L1 cache size. It was found in practice that the ideal values for $v$ are 32 on NVIDIA GPUs and 16 on Intel GPUs. Perhaps non-intuitively, the ideal value on AMD GPUs was seen to be v=32v=32.333These values are encoded at https://github.com/lammps/lammps/blob/4246fab5/src/KOKKOS/kokkos_type.h#L1379-L1387 Kokkos enables this explicit experimentation and tuning which is important for maximizing performance. A deeper discussion of the specific role of cache size can be found in section 4.4.
@@ 4.3.3. ComputeUi and ComputeDuidrj
Further optimizations were possible on the GPU. This is based on the critical observation that recursive polynomial evaluation is inherently compute bound. The only inputs are the base case and all components of $\bf u$ follow from that.
@@
This approach carries over directly to ComputeDuidrj except we now stage the matrices $\bf U$} and the derivative $\bf\vec{d}u$}. In place of the atomic accumulation into $\bf U$} we can inline the accumulation into the force $\bf\vec{F}$—which is a kernel fusion with ComputeDeidrj. We denote this fused kernel ComputeFusedDeidrj.
@@ 4.3.4. Work Batching and Instruction Level Parallelism
We have already seen work batching in ReaxFF when the the two sparse linear system solves were fused, described in section 4.2.3. In the sparse matrix-vector operation, the units of independent work are loading components of the right-hand-side vector, multiplying by the same matrix element, and accumulating into independent output vectors. Our limiter is loading the matrix elements. We can simultaneously hide the latency of loading both vectors behind the latency of loading the matrix elements. This is the origin of the performance improvement.
@@
We followed this same roadmap for implementing batching in the SNAP kernel. Our initial optimization efforts focused on NVIDIA H100. Limiters were identified using NVIDIA Nsight Compute. However, the idea of work batching to shift limiters and improve ILP is not specific to any one architecture. By introducing the batch size as a compile-time tunable parameter, we could find performance improvements on AMD MI300A with otherwise identical code.
@@
We list the performance improvements for all three kernels in table 2. As noted before, these optimizations were driven by a quantitative analysis of limiters on NVIDIA H100. We could then empirically re-tune the batching factors on AMD MI300A and see if we achieved a speed-up. In the “worst case”, a batching factor of $1$ would restore the original performance.
@@ 4.4. Cache Size Performance Impact
It can be nontrivial to undertake cross-architecture investigations of performance. The Kokkos library eases this challenge considerably by offering a single source of flexible code with which to do controlled comparisons. One application of this capability that we demonstrate is investigating the sizes of L1 cache and shared memory capacities, which have known differences between GPU architectures.
Generally speaking, GPUs have two sets of L1-like cache: a hardware and a software managed one. The latter is referred to as shared memory in CUDA, local data share in HIP, and local memory in SYCL.
The characteristics of these caches are one of the major differences of the GPU architectures used on exascale class systems (see table 1). The sizes of the L1 cache and shared memory are fixed on AMD MI250X and MI300A, as well as on Intel PVC, because they are discrete units. On the other hand, modern NVIDIA GPUs have a unified cache where the L1 and shared memory capacity can be dynamically shifted.
The ability to shift between L1 and shared memory on NVIDIA GPUs enables us to design an experiment to investigate the impact of L1 or shared memory size on kernel performance leveraging what CUDA calls the shared memory carveout. This is the percentage of the unified cache reserved for shared memory. With kernels that do not use shared memory but benefit from L1 cache, we expect performance to decrease with an increasing ratio of the combined cache being reserved for shared memory. For kernels which use shared memory and do not rely on automatic caching in L1 the opposite is expected, and for ones that leverage both L1 and shared memory the best performance should be achieved with some intermediate carveout value.
Indeed Kokkos has a built-in heuristic to optimize the carveout value depending on the characteristics of the kernel launched. For this study, we overwrote that heuristic and simply forced a specific carveout value to see its impact on four top kernels: the primary force computation kernel for the Lennard Jones potential PairComputeLJCut, and the three most time consuming kernels in SNAP, ComputeUi, ComputeYi, and ComputeFusedDeidrj. Kernel runtimes were measured using NVIDIA Nsight Systems. We also looked at the top four kernels for ReaxFF but did not see a significant dependence on the cache carveout (under 10%).
The effect of manually setting the cache carveout on NVIDIA H100-HBM3-SXM is shown in figure 3. The x-axis is the shared memory carveout and the y-axis is the performance of each kernel normalized by the performance from running at the default carveout.
The force kernel of Lennard Jones does not use shared memory, but does see significant benefits from large L1 cache sizes. At the maximum carveout for shared memory, which leaves only 32kB for L1, performance drops by about 50%; alternatively, increasing the L1 cache size from 32kB to 224kB increases performance by 85%. The ComputeYi kernel shows a similar behavior; in this case the difference comes from reuse of $\mathbf{U}_{j}$ matrices instead of neighbor atom coordinates.